Irene got a fridge for $300 while Andrew got one for $260. What is Andrew's percenage savings?
A)133.3%
B)1.3%
C)86.7%
D)13.3%

Answer by checkley71(8403)

(300-260)/300=40/300=.13333=13.33%
ANSWER D) IS CORRECT.